OpenStudy (anonymous):

How do I set up the following problem: f(x)=x3(to the third power) +1 with the points [2,3] and [-1,1] very new to calculus and pretty unsure of myself thank you!!

OpenStudy (anonymous):

f(x)=x^2+1 (2,3) (-1,1) what are you supposed to find?

OpenStudy (anonymous):

The problem reads: Find the average rate of change of the function over the given interval or intervals problem 1: f(x) =x^3+1 points: [2,3] [-1,1]

OpenStudy (dumbcow):

average rate of change is basically the slope formula if you drew a line from [2,3] to [-1,1] it would be tangent to some point on f(x) anyway: slope = (f(2) -f(-1))/(2-(-1))